40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Dec 11 09:31:47 2012, Ayaka, Update, LSC, Oplev update for improving sensitivity 7x
    Reply  Tue Dec 11 21:53:37 2012, Ayaka, Update, LSC, Oplev update for improving sensitivity OL_BSpit.pdfBSoplevservo.png
Message ID: 7812     Entry time: Tue Dec 11 21:53:37 2012     In reply to: 7808
Author: Ayaka 
Type: Update 
Category: LSC 
Subject: Oplev update for improving sensitivity 

[Rana, Ayaka]

The BS oplev pitch feedback came back.

OL_BSpit.pdf

The problem was that 300^2:0 filter was off. And I turned on all the low pass filters (ELP35), then the oplev servo does not seem to inject big noise into the arms as long as I see the spectra of POY and POX. These low-pass filters will be modified tomorrow so that the acoustic coupling noise is minimized.

BSoplevservo.png

ELOG V3.1.3-